# An Open-Source Vision-Language-Action Model

### v1

@@ 2 Related Work
A number of works have explored the use of VLMs for robotics, e.g., for visual state representations [12, 13], object detection [67], high-level planning [16], and for providing a feedback signal [68, 69, 70, 71]. Others integrate VLMs directly into end-to-end visuomotor manipulation policies [14, 15], but incorporate significant structure into the policy architecture or require calibrated cameras, which limits their applicability. A number of recent works have explored similar recipes to ours and directly fine-tuned large pretrained VLMs for predicting robot actions [7, 1, 72, 17, 18]. Such models are often referred to as vision-language-action models (VLAs), since they fuse robot control actions directly into VLM backbones. This has three key benefits: (1) it performs alignment of pretrained vision and language components on a large, Internet-scale vision-language dataset, (2) the use of a generic architecture, not custom-made for robot control, allows us to leverage the scalable infrastructure underlying modern VLM training [73, 74, 75] and scale to training billion-parameter policies with minimal code modifications, and (3) it provides a direct pathway for robotics to benefit from the rapid improvements in VLMs. Existing works on VLAs either focus on training and evaluating in single robot or simulated setups [72, 76] and thus lack generality, or are closed and do not support efficient fine-tuning to new robot setups [7, 1, 17, 18].
@@ 3.1 Preliminaries: Vision-Language Models
Neither SigLIP / DinoV2, nor Llama 2 release details about their training data, which likely consists of trillions of tokens of Internet-sourced image-text and text-only data respectively. The Prismatic VLM is fine-tuned on top of these components using the LLaVA 1.5 data mixture [43], which contains a total of approximately 1M image-text and text-only data samples from open-source datasets [79, 80, 29, 81, 42].
@@ 5.1 Direct Evaluations on Multiple Robot Platforms
Notably, OpenVLA performs comparably to RT-2-X on Google robot evaluations and significantly outperforms RT-2-X on BridgeData V2 evaluations despite being an order of magnitude smaller (7B vs. 55B parameters). Qualitatively, we find that both RT-2-X and OpenVLA exhibit markedly more robust behaviors than the other tested models, such as approaching the correct object when distractor objects are present, properly orienting the robot’s end-effector to align with the orientation of the target object, and even recovering from mistakes such as insecurely grasping objects (see https://openvla.github.io for qualitative rollout examples). RT-2-X achieves higher performance in semantic generalization tasks, as shown in Fig. 2, which is expected given that it uses larger-scale Internet pretraining data and is co-fine-tuned with both robot action data and Internet pretraining data to better preserve the pretraining knowledge, rather than being fine-tuned solely on robot data, like OpenVLA. However, OpenVLA performs comparably or better in all other task categories in both BridgeData V2 and Google robot evaluations. The performance difference can be attributed to a combination of factors: we curated a much larger training dataset for OpenVLA with 970k trajectories (vs. 350k for RT-2-X); we performed more careful cleaning of the training dataset and, e.g., filtered out all-zero actions in the Bridge dataset (see Appendix C for a detailed discussion); and OpenVLA uses a fused vision encoder that combines pretrained semantic and spatial features.
@@ 5.2 Data-Efficient Adaptation to New Robot Setups
While prior works mainly focused on directly evaluating VLAs “out-of-the-box” [16, 7, 1], effective fine-tuning of VLA models to new tasks and robot setups is largely unexplored, yet is key for their widespread adoption. In this section, we investigate OpenVLA’s ability to be quickly adapted to a new robot setup.
@@ 5.3 Parameter-Efficient Fine-Tuning
Table 1: Parameter-efficient fine-tuning evaluation. LoRA fine-tuning [26] achieves the best performance-compute trade-off, matching full fine-tuning performance while training only 1.4% of the model parameters. Mean success rate $\pm$ StdErr is computed across 33 rollouts per approach on select Franka-Tabletop tasks.
@@ 5.4 Memory-Efficient Inference via Quantization
Table 2: Performance with quantized inference. 4-bit quantization matches the performance of bfloat16 inference while reducing the GPU memory footprint by more than half. Mean success rates $\pm$ StdErr are computed across 8 representative BridgeData V2 tasks [6] and 80 rollouts per approach.
OpenVLA, a 7B-parameter model, consumes more memory at inference time than prior open-source generalist policies such as Octo, which has $<$100M parameters. We follow best-practices from LLM serving by saving and loading OpenVLA in bfloat16 precision for inference, which cuts the memory footprint in half, allowing us to serve OpenVLA on GPUs with only 16GB of GPU memory. In this section, we test whether we can further reduce the required memory for policy inference and broaden accessibility of VLA policies, by using modern quantization techniques developed for serving LLMs [86, 27]. These approaches load the weights of the network at lower precision, thereby trading off reduced memory requirements for potentially reduced inference speed and accuracy.
@@
As a result of the reduced inference speed, we observe a substantial performance decrease with 8-bit quantization: on the A5000 GPU we use for our evaluations, we can only run the model at 1.2Hz, which significantly changes the system dynamics compared to the training dataset for the 5Hz non-blocking controller used in the BridgeData V2 tasks. [^5] We attribute the performance loss mostly to low inference speed, since both 8-bit and 4-bit quantization achieve comparable token accuracy to bfloat16 inference when evaluated offline on training data. Notably, 4-bit quantization results in similar performance as bfloat16 half-precision inference despite requiring less than half the amount of GPU memory. 4-bit quantized models can run at 3Hz on the A5000, thus more closely matching the system dynamics during data collection.

### v2

@@ 2 Related Work
A number of works have explored the use of VLMs for robotics, e.g., for visual state representations [12, 13], object detection [67], high-level planning [16], and for providing a feedback signal [68, 69, 70, 71]. Others integrate VLMs directly into end-to-end visuomotor manipulation policies [14, 15], but incorporate significant structure into the policy architecture or require calibrated cameras, which limits their applicability. A number of recent works have explored similar recipes to ours and directly fine-tuned large pretrained VLMs for predicting robot actions [7, 1, 72, 73, 17, 18, 74]. Such models are often referred to as vision-language-action models (VLAs), since they fuse robot control actions directly into VLM backbones. This has three key benefits: (1) it performs alignment of pretrained vision and language components on a large, Internet-scale vision-language dataset, (2) the use of a generic architecture, not custom-made for robot control, allows us to leverage the scalable infrastructure underlying modern VLM training [75, 76, 77] and scale to training billion-parameter policies with minimal code modifications, and (3) it provides a direct pathway for robotics to benefit from the rapid improvements in VLMs. Existing works on VLAs either focus on training and evaluating in single robot or simulated setups [72, 73, 78, 74] and thus lack generality, or are closed and do not support efficient fine-tuning to new robot setups [7, 1, 17, 18].
@@ 3.1 Preliminaries: Vision-Language Models
SigLIP, DinoV2, and Llama 2 do not release details about their training data, which likely consists of trillions of tokens of Internet-sourced image-text, image-only, and text-only data respectively. The Prismatic VLM is fine-tuned on top of these components using the LLaVA 1.5 data mixture [43], which contains a total of approximately 1M image-text and text-only data samples from open-source datasets [81, 82, 29, 83, 42].
@@ 5.1 Direct Evaluations on Multiple Robot Platforms
Notably, OpenVLA performs comparably to RT-2-X on Google robot evaluations and significantly outperforms RT-2-X on BridgeData V2 evaluations despite being an order of magnitude smaller (7B vs. 55B parameters). Qualitatively, we find that both RT-2-X and OpenVLA exhibit markedly more robust behaviors than the other tested models, such as approaching the correct object when distractor objects are present, properly orienting the robot’s end-effector to align with the orientation of the target object, and even recovering from mistakes such as insecurely grasping objects (see https://openvla.github.io for qualitative rollout examples). RT-2-X achieves higher performance in semantic generalization tasks, as shown in Fig. 2, which is expected given that it uses larger-scale Internet pretraining data and is co-fine-tuned with both robot action data and Internet pretraining data to better preserve the pretraining knowledge, rather than being fine-tuned solely on robot data, like OpenVLA. However, OpenVLA performs comparably or better in all other task categories in both BridgeData V2 and Google robot evaluations. The performance difference can be attributed to a combination of factors: we curated a much larger training dataset for OpenVLA with 970k trajectories (vs. 350k for RT-2-X); we performed more careful cleaning of the training dataset and, e.g., filtered out all-zero actions in the Bridge dataset (see Appendix C for a detailed discussion); and OpenVLA uses a fused vision encoder that combines pretrained semantic and spatial features. See Appendix D for ablation analyses of these components.
@@ 5.2 Data-Efficient Adaptation to New Robot Setups
While prior works mainly focused on directly evaluating VLAs “out-of-the-box” [16, 7, 1], effective fine-tuning of VLA models to new tasks and robot setups is largely unexplored, yet is key for their widespread adoption. In this section, we investigate OpenVLA’s ability to be quickly adapted to a new real-world robot setup. (See Appendix E for fine-tuning experiments in simulation.)
@@ 5.3 Parameter-Efficient Fine-Tuning
Table 1: Parameter-efficient fine-tuning evaluation. LoRA fine-tuning achieves the best performance-compute trade-off, matching full fine-tuning performance while training only 1.4% of the model parameters. Mean success $\pm$ StdErr computed across 33 rollouts per approach on select Franka-Tabletop tasks (see Table 8 for details).
@@ 5.4 Memory-Efficient Inference via Quantization
Table 2: Performance with quantized inference. 4-bit quantization matches the performance of bfloat16 inference (our default approach) while reducing the GPU memory footprint by more than half. Mean success $\pm$ StdErr computed across 8 representative BridgeData V2 tasks [6] and 80 rollouts per approach (see Table 5 for details).
OpenVLA, a 7B-parameter model, consumes more memory at inference time than prior open-source generalist policies such as Octo, which has $<$100M parameters. We follow best-practices from LLM serving by saving and loading OpenVLA in bfloat16 precision for inference (our default approach), which cuts the memory footprint in half, allowing us to serve OpenVLA on GPUs with only 16GB of GPU memory. In this section, we test whether we can further reduce the required memory for policy inference and broaden accessibility of VLA policies, by using modern quantization techniques developed for serving LLMs [88, 27]. These approaches load the weights of the network at lower precision, thereby trading off reduced memory requirements for potentially reduced inference speed and accuracy.
@@
As a result of the reduced inference speed, we observe a substantial performance decrease with 8-bit quantization: on the A5000 GPU we use for our evaluations, we can only run the model at 1.2Hz, which significantly changes the system dynamics compared to the training dataset for the 5Hz non-blocking controller used in the BridgeData V2 tasks. [^5] We attribute the performance loss to low inference speed, since both 8-bit and 4-bit quantization achieve comparable token accuracy to bfloat16 inference when evaluated offline on training data. See Section D.4 for supporting details. Notably, 4-bit quantization results in similar performance as bfloat16 half-precision inference despite requiring less than half the amount of GPU memory. 4-bit quantized models can run at 3Hz on the A5000, thus more closely matching the system dynamics during data collection.
@@ B.1.3 Detailed BridgeData V2 Evaluation Results
Additionally, in Table 5, we provide the full evaluation results for the quantized inference experiments that were summarized in Table 2. For these evaluations, we test policies on 8 representative BridgeData V2 tasks spanning all task categories in the full evaluation suite.
Table 5: Full quantized inference results. Here we present the detailed version of the results shown in Table 2.
Category Task # Trials bfloat16 # Successes int8 # Successes int4 # Successes
Visual gen Put Eggplant into Pot (Easy Version) 10 9 7 9
Visual gen Put Eggplant into Pot 10 7 7 7
Visual gen Put Cup from Counter into Sink 10 5 3 7
Motion gen Lift Eggplant 10 6 4 7.5
Physical gen Put Carrot on Plate 10 6 5 7
Physical gen Lift AAA Battery 10 7 5 3
Semantic gen Take Purple Grapes out of Pot 10 8 8 9
Language grounding Put {Eggplant, Red Bottle} into Pot 10 9 7.5 8
Mean Success Rate 71.3 $\pm$ 4.8% 58.1 $\pm$ 5.1% 71.9 $\pm$ 4.7%
@@ B.2.1 Google Robot Evaluation Tasks
  12. 12.
    Move Coke Can to Taylor Swift (OOD): The robot is positioned in front of a platform with a can of Coke and photos of three different celebrities, including Taylor Swift. The robot’s goal is to grasp the can and move it to the photo of Taylor Swift. This task is OOD because the photos of the celebrities are unseen in the robot interaction data.
@@ B.3.2 Detailed Franka-Tabletop and Franka-DROID Evaluation Results
Additionally, in Table 8, we show the detailed version of the parameter-efficient fine-tuning experiment results summarized in Table 1. In these experiments, we use a representative subset of two Franka-Tabletop tasks, with both in-distribution and OOD variants: one narrow single-instruction task (“Put Carrot in Bowl”) and one diverse multi-instruction task (“Move <object> onto Plate”). We use the same number of training demonstrations used in Section 5.2 (50 and 150, respectively), which is delineated in Section B.3.1.
Table 8: Detailed parameter-efficient fine-tuning experiment results. Here we present the detailed task performance results summarized in Table 1.
# trials Full FT Last layer only Frozen vision Sandwich LoRA, r=32 LoRA, r=64
Franka-Tabletop (5Hz) “Put Carrot in Bowl” (in-distribution) 10 90.0 40.0 40.0 90.0 60.0 90.0
“Put Carrot in Bowl” (OOD) 5 40.0 0.0 40.0 0.0 60.0 40.0
“Move <object> onto Plate” (in-distribution) 12 79.2 33.3 50.0 75.0 75.0 62.5
“Move <object> onto Plate” (OOD) 6 41.7 33.3 58.3 41.7 75.0 66.7
Average 69.7$\pm$7.230.3$\pm$6.147.0$\pm$6.962.1$\pm$7.968.2$\pm$7.568.2$\pm$7.8%
@@
Appendix D Additional Experiments and Ablations
In this section, we conduct several additional experiments to analyze the effects of individual components of the OpenVLA model architecture and training scheme, as well as provide quantitative evidence for claims made in earlier sections of this work. We aim to answer the following questions:
  1. 1.
    How important is OpenX training and how does it impact OpenVLA’s performance (Section D.1)?
  2. 2.
    What effect does using a fused SigLIP-DinoV2 vision encoder have on OpenVLA’s performance, compared to using a SigLIP-only vision encoder (Section D.2)?
  3. 3.
    Is it better to fine-tune or freeze the vision encoder in OpenVLA (Section D.3)?
  4. 4.
    How do the quantized inference results discussed in Section 5.3 change when policy performance is disentangled from model inference speed (Section D.4)?
We discuss the experimental setup and results addressing each of the above questions sequentially in the following sections.
D.1 OpenX Training Data Ablation Experiments
As discussed in Section 3.3, OpenVLA is trained on a large dataset of robot embodiments, scenes, and tasks from the Open X-Embodiment dataset [1] (OpenX). In this section, we ablate the OpenX mixture and train a VLA policy solely on one robot dataset, to assess the impact of OpenX training on policy performance. Note that we have already observed the negative effect of ablating OpenX training in the fine-tuning regime, as discussed in Section 5.2 (see OpenVLA (Scratch)), but we discuss additional experiments on another robot embodiment in this section to provide more supporting evidence.
Experimental setup and tasks. We compare the original OpenVLA model with OpenVLA-Bridge, which is produced by taking the same pretrained VLM as OpenVLA (Prismatic VLM [44]) and fine-tuning it solely on BridgeData V2 [6] rather than the entire OpenX training mixture discussed in Appendix A. We evaluate OpenVLA and OpenVLA-Bridge on a subset of 8 representative tasks from the BridgeData V2 WidowX robot evaluation suite discussed in Section B.1.1. The tasks are listed in Table 9.
Results. Results for the OpenX training mixture ablation are shown in Table 9. By comparing OpenVLA with OpenVLA-Bridge, we see that performance drops drastically (reduction of 30 percent in absolute success rate), which demonstrates the importance of OpenX pretraining on final policy performance. Although the language grounding performance is not impacted, we observe performance reduction across all generalization categories. This result suggests that the large diversity of scenes, objects, and tasks in the OpenX training mixture is essential for unlocking improved generalization capabilities in the OpenVLA model.
Table 9: BridgeData V2 WidowX ablation experiment results. We evaluate various methods on a subset of 8 representative tasks to assess the importance of different components of the OpenVLA model architecture and training scheme. OpenVLA-Bridge is a version of OpenVLA without OpenX training (it is trained only on BridgeData V2), and OpenVLA-Bridge-SigLIP additionally ablates the fused vision backbone by removing the DinoV2 encoder (its vision backbone only consists of the SigLIP encoder). We observe that both OpenX training and the fused vision encoder improve policy performance, though the former has a much greater effect than the latter.
Category Task # Trials OpenVLA # Successes OpenVLA-Bridge # Successes OpenVLA-Bridge-SigLIP # Successes
Visual gen Put Eggplant into Pot (Easy Version) 10 10 8 8
Visual gen Put Eggplant into Pot 10 10 2 3
Visual gen Put Cup from Counter into Sink 10 7 4 2
Motion gen Lift Eggplant 10 7.5 5.5 6.5
Physical gen Put Carrot on Plate 10 8 4 1
Physical gen Lift AAA Battery 10 7 2 2
Semantic gen Take Purple Grapes out of Pot 10 4 3 3
Language grounding Put {Eggplant, Red Bottle} into Pot 10 7.5 8 7
Mean Success Rate 76.3 $\pm$ 4.8% 45.6 $\pm$ 5.6% 40.6 $\pm$ 5.5%
D.2 Dual vs. Single Vision Encoder Experiments
The OpenVLA model architecture consists of a fused vision backbone that combines the SigLIP [9] and DinoV2 [25] encoders. In this section, we ablate the DinoV2 component to assess the importance of using a dual vision encoder.
Experimental setup and tasks. We instantiate a model, OpenVLA-Bridge-SigLIP, which is a version of OpenVLA that is trained only on BridgeData V2 and consists of only the SigLIP encoder as the vision backbone. We compare this model with the OpenVLA-Bridge model discussed in the previous section (Section D.1), which shares the same model architecture as the original OpenVLA model and is only trained on Bridge robot data. Therefore, the only difference between OpenVLA-Bridge-SigLIP and OpenVLA-Bridge is that the former omits the DinoV2 encoder in the vision backbone. We evaluate these models on the same subset of 8 Bridge tasks described in the previous section.
Results. Results for the dual vision encoder ablation are shown in Table 9. The drop in performance from OpenVLA-Bridge to OpenVLA-Bridge-SigLIP implies that additionally including the DinoV2 encoder in the vision backbone improves policy performance. However, the 5 percent reduction in performance here is not as significant as the 30 percent drop in performance observed from ablating OpenX training. The low-level spatial features represented in DinoV2 appear to aid generalization in only some cases.
D.3 Fine-Tuned vs. Frozen Vision Encoder Experiments
As discussed in Section 3.4, prior work on VLMs observed higher performance from freezing the vision encoder than fine-tuning its parameters [44]. However, when training OpenVLA, we fine-tuned all 7B parameters in the model, including the SigLIP-DinoV2 vision backbone, as we discovered early on during development that fine-tuning the vision encoder led to higher-performing VLAs — a finding which held across various pretrained VLMs and model architectures. We discuss details of such findings below.
Experimental setup and tasks. In this section, we report the performance of two VLA policies produced by fine-tuning two different pretrained models from the Prismatic VLMs [44] repository on BridgeData V2. The two pretrained models are named SigLIP ViT-SO 224px and LLaVa v1.5 7B (Reproduction); see Karamcheti et al. [44] for details on their architectures and training mixtures. We evaluate both policies on various Bridge tasks shown in Table 10. Note that the evaluation configurations here differ from previously discussed Bridge evaluations, so the results are not directly comparable to results in other similar experiments.
Results. Results for the fine-tuned vs. frozen vision encoder experiments are shown in Table 10. We find that for both VLAs tested, fine-tuning the vision encoder leads to significantly higher success rates across various tasks. Qualitatively, in some cases, deploying the frozen vision encoder policies leads to unstable robot behaviors that are clearly suboptimal. Consequently, we decided early on during development to not conduct further experimentation with frozen vision encoders.
Table 10: Fine-tuned vs. frozen vision encoder experiment results. We evaluate the performance of fine-tuning (“Fine-Tuned”) vs. freezing the vision encoder (“Frozen Vision”) in two VLA policies built on top of two different pretrained VLMs from the Prismatic VLMs [44] repository. BridgeData V2 WidowX tasks shown here are performed in the same sink environment used for other Bridge experiments in this work (however, the initial environment configurations here differ, as these evaluations were conducted at an earlier stage in the project). We find that fine-tuning the vision encoder is crucial to obtain good policy performance. Certain frozen vision encoder evaluations were discontinued due to very poor (near-zero) performance and unstable robot behaviors. Among the evaluations where both frozen vision and fine-tuned approaches are tested, fine-tuning the vision encoder leads to 80.0% average success versus 46.7% average success from leaving it frozen.
SigLIP ViT-SO 224px LLaVa v1.5 7B (Reproduction)
Task # Trials Frozen Vision # Successes Fine-Tuned # Successes Frozen Vision # Successes Fine-Tuned # Successes
Put Eggplant into Pot 10 7 10 5 9
Put Corn on Plate 10 10 9 0 9
Mean Success Rate 85 95 25 90
Put { Eggplant, Red Bottle } into Pot 4 2 4 3
Put { Blue Cup, Pink Cup } on Plate 4 0 0 0
Lift { Cheese, Red Chili Pepper } 4 0 3 2
Put { Strawberry, Lime } into Pot 4 1 0 3
Move { Sushi, Grapes } 4 3 4 3
Mean Success Rate 30 55 55
D.4 Additional Quantized Inference Experiments: Disentangling Policy Performance and Model Inference Speed
In Section 5.3, we evaluated OpenVLA with different levels of precision at inference time: half precision (bfloat16), 8-bit quantization, and 4-bit quantization. 8-bit quantization led to lower BridgeData V2 performance relative to the other two approaches, and we hypothesized that the reduction in performance was caused by lower model inference speed from the operations used in 8-bit quantization. In this section, we conduct experiments to assess the veracity of this claim.
Specifically, we evaluate OpenVLA again with the three different levels of precision listed above, but now with blocking control. In other words, each action is fully executed on the robot before the next one is predicted by the policy and executed by the controller. This scheme controls system dynamics across methods with varying amounts of latency and thus allows us to test the quality of a policy’s action predictions, independent of its prediction speed. Effectively, the precision levels that have higher throughput – bfloat16 and 4-bit quantization – are forced to run slower to match the dynamics observed when deploying OpenVLA with 8-bit precision. Therefore, we expect OpenVLA’s performance with 8-bit precision to match the performance of bfloat16 and 4-bit precision under blocking control.
Experimental setup and tasks. We report the performance of OpenVLA with blocking control and quantized inference on the same subset of 8 BridgeData V2 tasks used in Section D.1 and Section D.2.
Results. Quantized inference experiment results with blocking control are shown in Table 11. Unlike in Table 2, where 8-bit quantization led to the worst rollout performance due to low inference speed, here we observe that 8-bit quantization performs comparably to bfloat16 precision and 4-bit quantization given that we evaluate with blocking control to remove the influence of varying inference speeds on task performance. This confirms our hypothesis about the effect of inference speed on 8-bit quantization performance in previous experiments (when using non-blocking control). We also see no substantial performance degradation when using the lowest precision, 4-bit, as also observed in Section 5.3.
Table 11: Quantized inference experiment results with blocking control. We report the success rate and standard error of OpenVLA on various BridgeData V2 WidowX tasks with bfloat16 precision (the default approach), 8-bit quantization (int8), and 4-bit quantization (int4) at inference time. All average success rates have overlapping error bars, which suggests that all methods perform comparably.
Category Task # Trials bfloat16 # Successes int8 # Successes int4 # Successes
Visual gen Put Eggplant into Pot (Easy Version) 10 10 10 10
Visual gen Put Eggplant into Pot 10 9 10 10
Visual gen Put Cup from Counter into Sink 10 5 5 3
Motion gen Lift Eggplant 10 8 7 7.5
Physical gen Put Carrot on Plate 10 10 10 10
Physical gen Lift AAA Battery 10 3 6 4
Semantic gen Take Purple Grapes out of Pot 10 2 2 2
Language grounding Put {Eggplant, Red Bottle} into Pot 10 9 9.5 8.5
Mean Success Rate 70.0 $\pm$ 5.1% 74.4 $\pm$ 4.9% 68.8 $\pm$ 5.2%
Appendix E LIBERO Simulation Experiments
Our previous discussions in Section 5.2 and Section 5.3 focused on adapting OpenVLA to novel real-world robot setups and tasks. This section explores adapting OpenVLA to simulated robot setups and tasks, specifically utilizing the LIBERO benchmark [116]. Our experimentation in simulation offers two key advantages:
  1. 1.
    Demonstration of versatility: We show that OpenVLA, despite having been pretrained exclusively on real-world robot data, can effectively adapt to simulated domains, overcoming potential disparities between real-world and simulated environments and dynamics.
  2. 2.
    Enhanced accessibility and reproducibility: Integration of OpenVLA into a publicly available simulation platform makes our model more accessible to other researchers, especially those who may not have access to robotic hardware. Additionally, simulated experiments are more easily reproduced than their real-world counterparts.
We discuss the experimental setup in Section E.1 and the results in Section E.2. We release the materials required to reproduce the experiments along with the OpenVLA codebase.
E.1 LIBERO Simulation Experimental Setup
Simulation setup and tasks. The LIBERO benchmark [116] consists of four task suites designed for studying lifelong learning in robotic manipulation, and the original paper therefore investigates both forward and backward transfer to a variety of tasks. In our experiments, we focus solely on supervised fine-tuning on the target task suite, measuring the performance of various policies trained via behavioral cloning on successful demonstrations of the tasks.
We perform experiments with the following four task suites, which each contain 10 tasks with 50 human-teleoperated demonstrations each:
  • LIBERO-Spatial consists of the same set of objects but different layouts, and tests the model’s understanding of spatial relationships.
  • LIBERO-Object consists of the same scene layouts but different objects, and tests the model’s understanding of object types.
  • LIBERO-Goal consists of the same objects and layouts but different task goals, and tests the model’s knowledge of different task-oriented behaviors.
  • LIBERO-Long (also called LIBERO-10) consists of long-horizon tasks with diverse objects, layouts, and tasks.
We make the following modifications to each of the training datasets above:
  1. 1.
    To accommodate methods requiring higher-resolution images (such as $256\times 256$px or $224\times 224$px), we regenerate all demonstrations at an increased resolution of $256\times 256$px. Originally, the dataset provided by the benchmark consists of $128\times 128$px images. We find that simply upscaling these images to $256\times 256$px results in poor image quality. Therefore, we choose to begin with higher-resolution images, which can be downscaled as necessary, ensuring higher image quality across various resolution requirements. These higher-resolution images were obtained by stepping through the simulation environments with the actions stored in the provided human-collected demonstrations and saving the images rendered by the simulator.
  2. 2.
    We filter out all “no-op” actions from the dataset, i.e., actions that have near-zero magnitude in the translation and rotation components and do not change the state of the robot’s gripper. We find that this simple data cleaning step is crucial for highly expressive single-step policies such as OpenVLA, which otherwise learn to imitate these no-op actions and consequently freeze indefinitely at certain states during evaluation.
  3. 3.
    We rotate all third-person images at both train and test time by 180 degrees because we observe that the LIBERO environments return images that are upside down on our hardware.
  4. 4.
    Since we train policies via imitation learning, which expects demonstrations to be successful, we replay all demonstrations in the corresponding simulation environments and filter out the demonstrations that fail to complete the task (as determined by the environments’ success criteria). As a result, we remove 68 of 500 LIBERO-Spatial demonstrations, 46 of 500 LIBERO-Object demonstrations, 72 of 500 LIBERO-Goal demonstrations, and 121 of 500 LIBERO-Long demonstratinos.
  5. 5.
    For all methods in our comparisons, we only utilize the static third-person camera images; we do not use the wrist camera images that are additionally provided in the original datasets. This is for sake of having fair comparisons, as OpenVLA’s visual inputs only consist of third-person camera images.
Comparisons. The methods that we compare include Diffusion Policy [^8] We use the implementation of Diffusion Policy that is described in the DROID dataset paper [11], which conditions action generation on DistilBERT [117] language embeddings of the task label. [3] trained from scratch, Octo [5] fine-tuned on the target dataset, and OpenVLA fine-tuned on the target dataset via LoRA ($r=32$) as described in Section 5.3. Each policy is trained independently on each of the task suites above (rather than training a single policy on all four suites combined). All policies are trained with the same set of demonstrations, so all methods benefit from the data cleaning steps described above.
Evaluation details. To ensure lower variance in the experimental results, all methods are evaluated across 500 trials for each task suite, and the reported performance is the average success rate over three random seeds (resulting in 1500 total trials per statistic). Although we modify the training datasets, as described earlier, we do not change the test environments but rather use the same initial environment configurations provided by the original LIBERO benchmark.
E.2 LIBERO Simulation Experimental Results
We present the LIBERO experimental results in Table 12. Importantly, we observe that OpenVLA can be effectively adapted to tasks in the LIBERO simulation environments, as it obtains highest average success rate and rank among the tested methods. However, we find that the overall margin between OpenVLA and the other methods are tighter here than in the real-world fine-tuning experiments discussed in Section 5.2. We attribute this to the fact that OpenVLA was pretrained with purely real-world robot data and no simulation data, which suggests that fine-tuning the model on simulated robot tasks may not be as effective as fine-tuning it on real-world tasks due to the domain gap between simulated and real-world environments and dynamics. We see evidence for this notion in the results obtained by Octo – another policy pretrained on large amounts of real-world robot data – which also only achieves a small boost in overall performance relative to a simple, strong baseline such as Diffusion Policy trained from scratch. We expect increased gains in performance for the pretrained and fine-tuned methods if simulation data is added to the pretraining data mixture.
Table 12: LIBERO simulation benchmark results. We report the success rate (SR) and standard error of each method for the four task suites in the LIBERO benchmark, averaged over three random seeds with 500 trials each. In addition, we show the ranking of each method within each task suite, where a rank of 1 indicates the strongest method in the suite and a rank of 3 indicates the weakest method. (The average ranking is important to note since it informs which method may be most suitable to use as a default for a variety of tasks; it is more informative than the average success rate, which is not normalized by individual task suite difficulty.) Overall, we find that fine-tuned OpenVLA achieves highest average success rate and rank, followed by fine-tuned Octo and then Diffusion Policy trained from scratch.
LIBERO-Spatial LIBERO-Object LIBERO-Goal LIBERO-Long Average
SR ($\uparrow$) Rank ($\downarrow$) SR ($\uparrow$) Rank ($\downarrow$) SR ($\uparrow$) Rank ($\downarrow$) SR ($\uparrow$) Rank ($\downarrow$) SR ($\uparrow$) Rank ($\downarrow$)
Diffusion Policy from scratch 78.3 $\pm$ 1.1% 3 92.5 $\pm$ 0.7% 1 68.3 $\pm$ 1.2% 3 50.5 $\pm$ 1.3% 3 72.4 $\pm$ 0.7% 2.5
Octo fine-tuned 78.9 $\pm$ 0.1% 2 85.7 $\pm$ 0.9% 3 84.6 $\pm$ 0.9% 1 51.1 $\pm$ 1.3% 2 75.1 $\pm$ 0.6% 2
OpenVLA fine-tuned (ours) 84.7 $\pm$ 0.9% 1 88.4 $\pm$ 0.8% 2 79.2 $\pm$ 1.0% 2 53.7 $\pm$ 1.3% 1 76.5 $\pm$ 0.6% 1.5